Fermi National Laboratory
 

Operation Summary for September 24, 2010 - September 27, 2010

 

The Integrated Luminosity for the period from 9/20/10 to 9/27/10 was 46.22 pb-1 . NuMI
reported receiving no protons on target during this same period due to target work.

Friday Sept 24

The day shift began with the Tevatron (TeV) off for LCW system repairs and sector accesses, with Operations monitoring the Recycler stash(272.6E10), with the Antiproton Source (Pbar) stacking (9.4E10), with the Linac, Booster, the and Main Injector (MI) operating normally, with the NuMI off for target work, with MiniBooNE and Meson FTBF T992 and T995 experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access.

During the day shift, the accesses lasted until 2 PM. MECAR went down while controls personnel tried to load it with new turn on orders; a local power cycle and reboot fixed the problem. Instrumentation personnel worked on the D1 BPMs.

During the evening shift, the TeV D1 BPM work finished up and Operations put the TeV into HEP shot setup at 6:58 PM. At 7:44 PM, the TeV quenched; no cause could be found. Operations had recovered the TeV cryo system and had the TeV in shot setup by the end of the shift.

Saturday Sept 25

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (232.8E10), with Pbar stacking (13.7E10), and with MiniBooNE taking beam.

During the midnight shift, operators halted shot setup when they found that the Tevatron Electron Lens (TEL2) solenoid was off. They reset it and put the TeV through a wet squeeze. Operations put the TeV into shot setup again and established store 8122 at 2:47 AM with an average initial luminosity (AIL) of 258E30. The TeV D1 BPMs stopped responding. Operations established beam to T992 and T995 at 6:26 AM.

During the day shift, Operations terminated store 8122 at 3:13 PM and put the TeV into shot setup.

During the evening shift, Operations established store 8123 at 4:28 AM with an AIL of 256.9E30. Controls issues with the Time Line Generator (TLG) caused several hours of downtime during the shift. Controls personnel investigated and reported at 10:56 PM that they had found and replaced a bad module.

Sunday Sept 26

The midnight shift began with Operations monitoring store 8123 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations established beam to T992 and T995 at 6:39 AM. Operations terminated store 8123 at 7:05 AM and then prepared D0 for an access into their collision hall.

During the day shift, D0 came out of access. Operations put the TeV into shot setup at 12:54 PM and established store 8126 at 2:10 PM with an AIL of 244E30. Operations reported that the TLG started hanging up again with the same symptoms as the day before. Controls personnel investigated. Operations reported that a Linac RF station (LRF3) started tripping repeatedly and would not remain on. Operators called in Linac personnel to investigate and they determined that the station’s driver tube needed to be replaced, which they began.

During the evening shift, Operations reported that the TLG and Linac RF work finished up at 9:48 PM. However, when Operations put Pbar into stacking mode, an Accumulator RF station (ARF1) would not come on. Operators investigated and found that a filament power supply was off and would not reset. Operators called in EE Support personnel and they had the supply online by 10:56 PM.

Monday Sept 27

The midnight shift began with Operations monitoring store 8126 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8126 at 4:55 AM, put the TeV into shot setup, and established store 8127 at 6:21 AM with an AIL of 251.54E30. Operations established beam to T992 and T995 at 7:15 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac personnel to work on LRF3. NuMI reported that they must replace their target; the new target won’t be ready for a month.

 

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/27/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory